### v1

@@ 4.1.5 Structure in Control Predictions
[…]
For $\alpha=0.15$, the model correctly captures the data’s location in the three-dimensional space. The hypersurfaces are barely visible due to a large amount of noisy data in between, which is expected since the training data for lower thrust levels is also noisy, and the ACT data for $\alpha=0.15$ shows only a weak hypersurface structure. For $\alpha=0.55$, the model accurately predicts the solution structure. It also correctly predicts the shooting time $\tau_{s}$, which changes significantly across different thrust levels. Notably, the model successfully reproduces the trend where most of the data between the hypersurfaces exhibits longer shooting times compared to the rest of the data.
@@
Figure 13: Comparison of the structure in the costate space, with 5,000 data points per method.
@@ 4.1.6 Model Performance
Table 4: Diffusion model sampling results for 300,000 initializations per thrust level, compared to samples from the ACT method. Solutions/minute include both sampling and solving times.
$\alpha=0.15$ $\alpha=0.55$ $\alpha=0.95$
@@
[…]
Sampling from the diffusion model enables solution generation at a high rate for the test problems. Although the runtime is longer compared to the ACT method, and additional GPU sampling time must be accounted for, the number of solutions generated per minute is significantly higher with diffusion model samples. For both methods, runtime decreases as the thrust level increases, which is due to shorter shooting times and faster trajectory propagation in the preliminary screening algorithm. The higher runtimes for diffusion model samples compared to the ACT method are a result of the greater number of feasible solutions found, which requires additional operations to save the results. This time could be further reduced by optimizing the code for these operations and only saving the essential parameters of each solution. The performance improvement is most notable for $\alpha=0.95$, with a factor of more than 75, while for the more challenging problem of $\alpha=0.15$, the solutions per minute still increase by a factor of over 20. The solutions generated by the diffusion model are of high quality with respect to the objective, as evidenced by a low average $\Delta v$. Since the Pareto front was explicitly targeted during the training data generation, the solutions produced by the diffusion model are also close to Pareto optimality.
@@ 4.2.1 Problem Description
Initial state $[\boldsymbol{r}_{0}^{T},\boldsymbol{v}_{0}^{T}]$ [NU] $[0.0763,-0.2786,-0.0445,0.9971,0.7897,-0.2632]$
Specific energy initial state [NU] $-2.5517$
Terminal state $[\boldsymbol{r}_{f}^{T},\boldsymbol{v}_{f}^{T}]$ [NU] $[0.8380,0.0579,0.0110,0.0427,0.0464,-0.0507]$
Specific energy halo orbit [NU] $-1.5842$
@@
The mission optimized in this problem is similar to the problem solved by Li et al. [2] and involves a transfer from an Earth orbit to a halo orbit around libration point $\mathcal{L}_{1}$. An example trajectory is shown in Fig. 14. Starting from the initial geostationary transfer orbit (GTO), the orbital energy is increased through a fixed-time tangential low-thrust spiral. The endpoint of this spiral serves as the initial boundary condition for the trajectory optimization problem. The dynamical structures of the CR3BP system are actively incorporated into the trajectory design by targeting the halo orbit through a final coast, $\tau_{f}$, along its stable interior invariant manifold. The terminal boundary condition is defined as any point on a fixed arc of this manifold, within the coasting bounds marked in Fig. 14. The specific point on the manifold arc that serves as the boundary condition is determined by the preliminary screening algorithm and parameterized through the final coast time. The detailed problem parameters are provided in Table LABEL:tab:problem_parameters_GTO_Halo. The bounds on the final coast time correspond to the time required to coast from the coasting bounds, marked in Fig. 14, to the target position on the halo orbit. The target position is the seed point of an arbitrarily chosen manifold arc on the halo orbit with a specified orbital energy. The spacecraft parameters are consistent with those used in Li et al. [2]. The maximum thrust, $T_{max}$, is treated as a variable problem parameter and is varied within a specified range. The constants used to normalize the distance, time, and mass units to the system of natural units in this section are also listed in Table LABEL:tab:problem_parameters_GTO_Halo. Unless otherwise specified, all variables in this section are normalized using these natural units. Solving this problem presents a greater challenge than the Europa DRO transfer described earlier, as it involves a three-dimensional transfer with a wide range of qualitatively different solutions and a dominant forcing component provided by the primary.
@@ 4.2.2 Data Generation
Convergence tolerance [NU] $0.001$ SNOPT mode optimal
$\varphi_{0}$ [rad] [0.1, 1.25]0.11.25[-0.1,\,1.25][ - 0.1 , 1.25 ] $\dot{\varphi}_{0}$ [rad/TU] [15.0, 0.0]15.00.0[-15.0,\,0.0][ - 15.0 , 0.0 ]
$\beta_{0}$ [0.4, 1.3]0.41.3[0.4,\,1.3][ 0.4 , 1.3 ] $\dot{\beta}_{0}$ $[-6.0,\,-1.0]$
$S_{0}-\lambda_{m,0}m_{0}/c$ [NU] [0.08, 0.4]0.080.4[0.08,\,0.4][ 0.08 , 0.4 ] $\dot{S}_{0}$ [NU] [0.5, 0.9]0.50.9[-0.5,\,0.9][ - 0.5 , 0.9 ]
@@ 4.2.3 Solution Structure
The structure of the six-dimensional costate space is examined by applying a Principal Component Analysis (PCA) to the training dataset. The data is projected onto the space defined by the first three principal components, labeled $PC1$, $PC2$, and $PC3$. The explained variance ratios for these components are $59.98\,\%$ for $PC1$, $27.01\,\%$ for $PC2$, and $12.50\,\%$ for $PC3$. Together, these three principal components account for over $99\,\%$ of the variance in the data, effectively capturing the vast majority of the information. The costate vectors are mapped to this new three-dimensional space using the transformation:
@@ 4.2.4 Structure in Control Predictions
Figure 17: Comparison of the structure in the costate space, with 5,000 data points per method.
@@ 4.2.5 Model Performance
Figure 18: Trajectories corresponding to control vectors from different diffusion timesteps $n$ for the GTO halo transfer with $\alpha=0.45$.
The process by which the diffusion model generates a high-quality initial guess for the costate vector, which allows to quickly solve for a feasible trajectory is visualized in Fig. 18. The model begins with noisy data sampled from a multivariate standard normal distribution and iteratively reconstructs structure in the data through a series of timesteps guided by the pre-trained neural network. Each trajectory in Fig. 18, corresponds to the trajectory resulting from the prediction for $\boldsymbol{z}_{0}$ after $(N-n)$ diffusion timesteps. The displayed trajectories are generated by propagating the initial state and initial costate guess forward in time. For consistency, the shooting time $\tau_{s}$ and final coast time $\tau_{f}$ for all displayed trajectories are set according to the values from the final converged trajectory. The first guess at $n=1$,000 results in an unreasonable trajectory that spirals into the Earth. As $n$ decreases, the model’s predictions become more reasonable and the magnitude of the constraint violations at the trajectory endpoint decreases. The actual sample at $n=0$ provides a good approximation of the solution. This guess is then used to warm-start the numerical solver SNOPT, which requires less than $10\,\mathrm{s}$ to converge to a feasible trajectory.
Although this process is visualized for just one example case, the overall high quality of diffusion model samples is confirmed across the testing thrust levels $\alpha=0.05$, $\alpha=0.45$ and $\alpha=0.95$, by comparing it to initial guesses sampled from a uniform distribution (referred to as uniformly sampled - US) and samples from the ACT method in Table 7. For each method, the results are based on 100,000 samples per thrust level, which are first filtered using the preliminary screening algorithm with a global search tolerance of $\delta=0.05$ and then solved with a maximum SNOPT runtime of $30\,$\mathrm{s}$$.
@@
Table 7: Comparison of solver warm-start performance for 100,000 initializations per method. Solutions per minute include both solving and sampling time for the diffusion model.
$\alpha=0.05$ $\alpha=0.45$ $\alpha=0.95$
@@ 5 Conclusion
A main challenge of this method is the large amount of data required to train the model, leading to a big upfront cost for data generation. Since the initial guesses sampled from the model are generally only as good as the data used to train it, this framework still needs to be combined with other methods which can generate high quality solutions. In future work, the framework will be further improved by generating training data at smaller intervals of the thrust level, with fewer data points per thrust level. Additionally, the training cost will be reduced by training the model on principal components, which capture the majority of information in the data. Furthermore, the application of the framework will be extended by conditioning it on different problem parameters and multiple parameters simultaneously. Examples of such conditional problem parameters could be the boundary conditions, represented as energy levels of the initial or target orbit, the mass parameter $\mu$ of the CR3BP system or the objective value itself. These extensions will further prepare the framework for potential application during preliminary mission design and potential enhancement of autonomous GNC capabilities.

### v2

@@ 4.1.5 Structure in Control Predictions
The diffusion model’s ability to predict changes in the structure of costate variables is demonstrated by testing it on new thrust levels that were not included in the dataset. We compare the generated costate samples from the model to separately generated solutions using our global search strategy. This comparison is shown for the thrust levels $\alpha=0.15$, $\alpha=0.55$, and $\alpha=0.95$ in Fig. 13. In line with the plots shown in Fig. 9, the costate samples are displayed in the 3D space spanned by $\lambda_{r0,2}$, $\lambda_{v0,1}$ and the transformed costate variable $\lambda^{\prime}$ from Eq. (32). Each scatter point corresponds to an initial costate that, when propagated forward for the shooting time indicated by the color plot, generates a candidate trajectory. In addition to the data directly sampled from the model (which is not necessarily feasible), the filtered diffusion model data containing only feasible solutions is presented, as well as separately generated feasible solutions based on ACT initializations. This reference data is obtained by applying the preliminary screening algorithm after sampling from the model, filtering the solutions based on feasibility with a tolerance of $\delta=$1\text{\times}{10}^{-4}$\$.
For $\alpha=0.15$, the model correctly captures the data’s location in the three-dimensional space. The hypersurfaces are barely visible due to a large amount of noisy data in between, which is expected since the training data for lower thrust levels is also noisy, and the ACT data for $\alpha=0.15$ shows only a weak hypersurface structure. For $\alpha=0.55$, the model accurately predicts the solution structure. It also correctly predicts the shooting time $\tau_{s}$, which changes significantly across different thrust levels. Notably, the model successfully reproduces the trend where most of the data between the hypersurfaces exhibits longer shooting times compared to the rest of the data. This is likely due to the solutions with longer shooting times being further away from Pareto optimality, therefore not matching the structure of most of the other solutions, which are close to the Pareto front.
@@
Figure 13: Comparison of the structure in the costate space, with 5,000 data points per method for $\alpha=0.15\,(T_{max}=0.75\,\mathrm{N})$, $\alpha=0.55\,(T_{max}=2.74\,\mathrm{N})$ and $\alpha=0.95\,(T_{max}=4.73\,\mathrm{N})$.
@@ 4.1.6 Model Performance
Table 4: Diffusion model sampling results compared to samples from the ACT method for 300,000 initializations per thrust level and a feasibility tolerance of $10^{-4}$. Solutions/minute include both sampling and solving times.
$\alpha=0.15\,(T_{max}=0.75\,\mathrm{N})$ $\alpha=0.55\,(T_{max}=2.74\,\mathrm{N})$ $\alpha=0.95\,(T_{max}=4.73\,\mathrm{N})$
@@
For this study, the preliminary screening algorithm was used with a convergence tolerance of $\delta=$1\text{\times}{10}^{-4}$\$, corresponding to the global search strategy represented by the lower path in Fig. 2. For the ACT method, the feasibility ratio remains close to $0.0\,\%$ across all three thrust values. Despite this, solutions are generated at a high rate of more than one per minute due to the low runtime of the ACT method. This highlights the strength of the preliminary screening algorithm which evaluates the feasibility of a solution in only a fraction of a second, allowing thousands of solutions to be processed in a short period. Due to the high quality of the diffusion model samples, the feasibility ratios are increased by more than two orders of magnitude. This is remarkable, considering that the model itself has no direct knowledge of the dynamics and constraints of the transfer. The feasibility ratio for $\alpha=0.95$ is lower than the values for $\alpha=0.15$ and $\alpha=0.55$. This aligns with the results from the previous section, where the costate structure for $\alpha=0.95$ was not captured as accurately. Nonetheless, the feasibility ratio remains significantly higher compared to samples from the ACT method.
Sampling from the diffusion model enables solution generation at a high rate for the test problems. Although the runtime is longer compared to the ACT method, and additional GPU sampling time must be accounted for, the number of solutions generated per minute is significantly higher with diffusion model samples. For both methods, runtime decreases as the thrust level increases, which is due to shorter shooting times and faster trajectory propagation in the preliminary screening algorithm. The higher runtimes for diffusion model samples compared to the ACT method are a result of the greater number of feasible solutions found, which requires additional operations to save the results. This time could be further reduced by optimizing the code for these operations and only saving the essential parameters of each solution. The performance improvement is most notable for $\alpha=0.95$, with a factor of more than 75, while for the more challenging problem of $\alpha=0.15$, the solutions per minute still increase by a factor of over 20. The solutions generated by the diffusion model are of high quality with respect to the objective, as evidenced by a low average $\Delta v$. Since the Pareto front was explicitly targeted during the training data generation, the solutions produced by the diffusion model are also close to Pareto optimality. This is further illustrated in Fig. 14, where the Pareto front for the diffusion model solutions is visualized.
Figure 14: A subset of 5,000 solutions generated from diffusion model samples for the three testing thrust levels, displayed in the $\Delta v$-$\tau_{s}$ plane. A pareto front is uncovered in all three cases.
@@ 4.2.1 Problem Description
Initial state $\boldsymbol{r}_{0}^{T},\boldsymbol{v}_{0}^{T}] [$NU] [0.0763,0.2786,0.0445,0.9971,0.7897,0.2632][0.0763,-0.2786,-0.0445,0.9971,0.7897,-0.2632]
Specific energy initial state [NU] 2.5517-2.5517
Terminal state $\boldsymbol{r}_{f}^{T},\boldsymbol{v}_{f}^{T}] [$NU] [0.8380,0.0579,0.0110,0.0427,0.0464,0.0507][0.8380,0.0579,0.0110,0.0427,0.0464,-0.0507]
Specific energy halo orbit [NU] 1.5842-1.5842
@@
The mission optimized in this problem is similar to the problem solved by Li et al. [2] and involves a transfer from an Earth orbit to a halo orbit around libration point $\mathcal{L}_{1}$. An example trajectory is shown in Fig. 15. Starting from the initial geostationary transfer orbit (GTO), the orbital energy is increased through a fixed-time tangential low-thrust spiral, which is only computed once and is identical for all solutions. The endpoint of this spiral serves as the initial boundary condition for the trajectory optimization problem. The dynamical structures of the CR3BP system are actively incorporated into the trajectory design by targeting the halo orbit through a final coast, $\tau_{f}$, along its stable interior invariant manifold. The terminal boundary condition is defined as any point on a fixed arc of this manifold, within the coasting bounds marked in Fig. 15. The specific point on the manifold arc that serves as the boundary condition is determined by the preliminary screening algorithm and parameterized through the final coast time. The detailed problem parameters are provided in Table LABEL:tab:problem_parameters_GTO_Halo. The bounds on the final coast time correspond to the time required to coast from the coasting bounds, marked in Fig. 15, to the target position on the halo orbit. The target position is the seed point of an arbitrarily chosen manifold arc on the halo orbit with a specified orbital energy. The spacecraft parameters are consistent with those used in Li et al. [2]. The maximum thrust, $T_{max}$, is treated as the conditional problem parameter and is varied within a specified range. The constants used to normalize the distance, time, and mass units to the system of natural units in this section are also listed in Table LABEL:tab:problem_parameters_GTO_Halo. Unless otherwise specified, all variables in this section are normalized using these natural units. Solving this problem presents a greater challenge than the Europa DRO transfer described earlier, as it involves a three-dimensional transfer with a wide range of qualitatively different solutions and a dominant forcing component provided by the primary.
@@ 4.2.2 Data Generation
Feasibility tolerance [NU] $0.001$ SNOPT mode optimal
$\varphi_{0}$ [rad] [0.1, 1.25][-0.1,\,1.25] $\dot{\varphi}_{0}$ [rad/TU] [15.0, 0.0][-15.0,\,0.0]
$\beta_{0}$ [0.4, 1.3][0.4,\,1.3] $\dot{\beta}_{0}$ [6.0,1.0][-6.0,\,-1.0]
$S_{0}-\lambda_{m,0}m_{0}/c$ [NU] [0.08, 0.4][0.08,\,0.4] $\dot{S}_{0}$ [NU] [0.5, 0.9][-0.5,\,0.9]
@@ 4.2.3 Solution Structure
The structure of the six-dimensional costate space is examined by applying a Principal Component Analysis (PCA) to the training dataset. PCA is a a linear dimensionality-reduction method that orthogonally transforms the data into principal components ranked by the variance they explain [46]. We project the data onto the space defined by the first three principal components, labeled $PC1$, $PC2$, and $PC3$. The explained variance ratios for these components are $59.98\,\%$ for $PC1$, $27.01\,\%$ for $PC2$, and $12.50\,\%$ for $PC3$. Together, these three principal components account for over $99\,\%$ of the variance in the data, effectively capturing the vast majority of the information. The costate vectors are mapped to this new three-dimensional space using the transformation:
@@ 4.2.4 Structure in Control Predictions
Figure 18: Comparison of the structure in the costate space, with 5,000 data points per method. Results are shown for thrust levels $\alpha=0.05\,(T_{max}=0.525\,\mathrm{N})$, $\alpha=0.45\,(T_{max}=0.725\,\mathrm{N})$ and $\alpha=0.95\,(T_{max}=0.975\,\mathrm{N})$.
@@ 4.2.5 Model Performance
Figure 19: Trajectories corresponding to control vectors from different diffusion timesteps $n$ for the GTO halo transfer with $\alpha=0.45\,(T_{max}=0.725\,\mathrm{N})$.
The process by which the diffusion model generates a high-quality initial guess for the costate vector, which allows to quickly solve for a feasible trajectory is visualized in Fig. 19. The model begins with noisy data sampled from a multivariate standard normal distribution and iteratively reconstructs structure in the data through a series of timesteps guided by the pre-trained neural network. Each trajectory in Fig. 19, corresponds to the trajectory resulting from the prediction for $\boldsymbol{z}_{0}$ after (Nn)(N-n) diffusion timesteps. To stay consistent with the reverse diffusion process visualized in Fig. 5, the timestep label decreases from $n=N$ to $n=0$. The displayed trajectories are generated by propagating the initial state and initial costate guess forward in time. For consistency, the shooting time $\tau_{s}$ and final coast time $\tau_{f}$ for all displayed trajectories are set according to the values from the final converged trajectory. The first guess at n=1n=1,000 results in an unreasonable trajectory that spirals into the Earth. As $n$ decreases, the model’s predictions become more reasonable and the magnitude of the constraint violations at the trajectory endpoint decreases. The actual sample at $n=0$ provides a good approximation of the solution. This guess is then used to warm-start the numerical solver SNOPT, which requires less than $10\,\mathrm{s}$ to converge to a feasible trajectory.
Although this process is visualized for just one example case, the overall high quality of diffusion model samples is confirmed across the testing thrust levels $\alpha=0.05$, $\alpha=0.45$ and $\alpha=0.95$, by comparing it to initial guesses sampled from a uniform distribution (referred to as uniformly sampled - US) and samples from the ACT method in Table 7. For each method, the results are based on 100,000 samples per thrust level, which are first filtered using the preliminary screening algorithm with a global search tolerance of $\delta=0.05$ and then solved with a maximum SNOPT runtime of $30\,$\mathrm{s}$\$ and a feasibility tolerance of $10^{-3}$. We perform these tests with the same coarse tolerances used for the training data and provide a study including tighter tolerances in the subsequent section.
@@
Table 7: Comparison of solver warm-start performance for 100,000 initializations per method and a feasibility tolerance of $10^{-3}$. Solutions per minute include both solving and sampling time for the diffusion model.
$\alpha=0.05\,(T_{max}=0.525\,\mathrm{N})$ $\alpha=0.45\,(T_{max}=0.725\,\mathrm{N})$ $\alpha=0.95\,(T_{max}=0.975\,\mathrm{N})$
@@
4.2.6 Tolerance Refinement
A necessary property of any useful solutions found from a global search using low-fidelity physics models, as is done in this paper, is that the solutions can be continued into high-fidelity models with tight feasibility tolerances. In the global search phase, where computational speed is important, the feasibility tolerances can be looser as long as they can be tightened at a later time. The feasibility tolerances used in this study ($10^{-4}$ NU for the Europa DRO transfer and $10^{-3}$ NU for the GTO Halo transfer) are coarse by mission‐design standards. As the aim of the paper is to demonstrate the ability to learn the global control structure using generative machine learning given a subset of data solved at a specified tolerance, we should expect that the trained model will provide control predictions that satisfy this specified tolerance and may struggle to provide as many predictions that satisfy a tighter tolerance.
[…]
To reiterate, the aim of this paper is to show that our framework learns the costate distribution of the fully parameterized problem, including inaccuracies that arise from the problem setup, control transcription, and numerical solver [3]. Because the training data was generated with relatively coarse solver tolerances, the diffusion model samples match that distribution. When tighter tolerances are imposed, some samples fail to converge because the new parameterization defines a slightly different distribution. If the diffusion model was trained on data computed with a $10^{-6}$ NU tolerance, it should yield samples that converge more often under these tight tolerances. Generating data for training under the tighter tolerances will be more computationally expensive and our ongoing work investigates fine-tuning the diffusion model with a Markov Chain Monte Carlo algorithm, which may allow the generation of samples that satisfy tighter tolerances with limited new training data from the tight tolerance case.
Table 8: Percentage of diffusion model solutions for the GTO Halo transfer ($\alpha=0.95$) that remain feasible when warm-starting SNOPT under increasingly tighter feasibility tolerances with a maximum solving time of $30\,\mathrm{s}$. Each result is based on 1,0001{,}000 initial solutions for a feasibility tolerance of $10^{-3}$.
Feasibility Tolerance [NU] $10^{-3}$ $10^{-4}$ $10^{-5}$ $10^{-6}$
Convergence Rate [%] 100.0 76.1 74.6 68.1
@@ 5 Conclusion
A main challenge of this method is the large amount of data required to train the model, leading to a big upfront cost for data generation. Since the initial guesses sampled from the model are generally only as good as the data used to train it, this framework still needs to be combined with other methods which can generate high quality solutions. In future work, the framework will be further improved by generating training data at smaller intervals of the thrust level, with fewer data points per thrust level. Additionally, the training cost will be reduced by training the model on principal components, which capture the majority of information in the data. Furthermore, the application of the framework will be extended by conditioning it on different problem parameters and multiple parameters simultaneously. Examples of such conditional problem parameters could be the boundary conditions, represented as energy levels of the initial or target orbit, the mass parameter $\mu$ of the CR3BP system or the objective value itself. These extensions will further prepare the framework for potential application during preliminary mission design and, when combined with tighter tolerances and a higher-fidelity model, for potential autonomous on-board applications.
@@ 7 Conflict of Interest
B. Qualitative study of different diffusion model parameters
Figure 20: Structure in the costate space for diffusion-model samples with varying hyper-parameters ($\alpha=1.0$, Europa DRO transfer). The reference model (parameters in Table 3) is shown in the first panel; each remaining panel varies only the indicated parameter. Each plot depicts a subset of 5 0005\,000 samples.